# Planet as a Brain: Towards Internet of AgentSites based on AIOS Server

### v1

@@ Abstract
In this paper, we introduce AIOS Server, a runtime framework designed to host agents and support global-scale collaboration among decentralized agents. AIOS Server provides a standardized communication protocol leveraging the Model Context Protocol (MCP) and JSON-RPC to enable structured agent-agent or human-agent interactions. Each AIOS node operates as an independent server, capable of hosting and executing agents, while supporting peer-to-peer coordination without reliance on centralized orchestration.
Based on AIOS Server, we further present the world’s first practically deployed Internet of Agentsites (AIOS-IoA), including AgentHub for agent registration and management as well as AgentChat for interactive communication, at https://planet.aios.foundation. Furthermore, we prototype an agent discovery mechanism based on Distributed Hash Tables (DHT) and a Gossip protocol, which serves as the search engine for the internet of agentsites, enabling scalable and resilient agent registry and lookup on this new internet.
Our evaluation demonstrates that AIOS Server achieves low-latency communication, efficient task delegation, and robust coordination in agent networks. This work provides a practical foundation for building the Internet of Agentsites — a new paradigm where autonomous agents become first-class citizens of the web. The implementation is available at https://github.com/agiresearch/AIOS.Server and will be integrated into the AIOS main branch at https://github.com/agiresearch/AIOS.
@@ 1 Introduction
The rapid progress of large language models (LLMs) has led to the emergence of autonomous agents capable of planning, reasoning, and interacting with humans and other agents in structured environments [17, 22, 21, 18]. These agents increasingly act beyond static response generation, demonstrating memory, tool usage, and long-term task execution abilities. However, most existing agent-based systems remain confined within centralized platforms, limiting openness, interoperability, and scalability in multi-agent ecosystems.
@@
This work builds on recent advances in agent system infrastructures [12, 7, 4], memory-enhanced LLM agents [21, 18], and agent communication protocols [1, 5, 11]. AIOS Server emphasizes modularity, interoperability, and robustness as foundational elements for open and scalable agent ecosystems. This work contributes a practical foundation for building the next-generation Internet of Agentsites, where autonomous agents operate as first-class citizens of the web, capable of decentralized collaboration, search, and interaction.
Recent works such as [3] explored the idea of the internet of agents as a multi-agent framework in terms of agent team up and collaboration. However, our work is different in that we do not focus on agent team up and collaboration, but on the physical, decentralized and real-world deployment of agents across the globe (as shown in Figure 1), just like the physical Internet of Websites, where each website runs on a server at a certain location on the planet, and servers are connected through the internet to constitute the world wide web (WWW). As a result, our infrastructure is denoted as the Internet of Agentsites for differentiation and for highlighting the analogy with the Internet of Websites.
@@ 2.1 LLM-Powered Agent Systems
Recent advances in large language models (LLMs) have enabled the creation of autonomous agents that can interpret tasks, use tools, and maintain state over long interactions [17, 22, 25, 23]. These agents are no longer limited to single-turn responses—they can reason, remember, and act in structured workflows [21, 18]. Frameworks such as Camel [9], OpenAGI [6], AIOS [12], AutoGen [19] and MetaGPT [7] define agent execution workflows and pipelines, where agents assume roles to work on task decomposition and decision-making either individually or collaboratively. Memory-augmented systems, including A-Mem and Workflow Memory [21, 18], provide persistent state tracking, improving contextual understanding and coherence. LLM-based agents now actively support applications in software development [15], web interaction [8, 4], and interactive simulations [16], extending their practical utility across domains.
@@ 2.2 Agent Communication and Protocol Design
Effective coordination in agent systems requires clear, interpretable communication. Structured interaction protocols—such as ReAct [22] and intent-based dialogue models—enable agents to plan, reason, and act based on conversational or environmental context. Recent works explore debate and discussion as communication primitives to improve factuality and reasoning depth [5, 11, 3, 14, 19]. Message-oriented protocols like JSON-RPC are widely adopted to support structured, machine-readable interaction between agent components or subsystems. Emerging platforms such as AgentVerse [2], OS-Copilot [20], and Formal-LLM [10] emphasize modular communication layers, enabling agents to interact with tools, APIs, and other services in well-defined formats.
@@ 2.3 Infrastructure and Decentralized Execution
Traditional AI agents often depend on centralized backends for registration, execution, and orchestration. This architecture limits scalability, introduces single points of failure, and reduces system adaptability. Recent efforts propose decentralized execution environments that distribute agent hosting and discovery across independent nodes. Platforms like AIOS (AI Agent Operating System) [12] highlight the need for developer-friendly, general-purpose agent development, hosting and execution frameworks. Our system builds on this foundation by introducing a runtime that integrates agent communication, registration, and discovery into a fully distributed infrastructure. Security and evaluation remain key concerns. Benchmarking frameworks such as ASB [24] and GAIA [13] provide standardized settings for assessing agent robustness, safety, and coordination capabilities at scale.
@@ 6 Experiments
We evaluate the AIOS server communication framework under local and cloud-based deployments. Experiments assess three core metrics: latency, throughput, and agent registration efficiency in a decentralized setting.
@@ 6.4.1 Global Node Distribution
Refer to caption (b) Eurasian view showing nodes in London, Singapore, and Tokyo.

### v3

# {textblock}

@@ Abstract
In this paper, we introduce AIOS Server, a runtime framework designed to host agents and support large-scale collaboration among decentralized agents. AIOS Server provides a standardized communication protocol leveraging the Model Context Protocol (MCP) and JSON-RPC to enable structured agent-agent or human-agent interactions. Each AIOS node operates as an independent server, capable of hosting and executing agents, while supporting peer-to-peer coordination without reliance on centralized orchestration.
Based on the AIOS Server, we further present the world’s first practically deployed Internet of Agentsites (AIOS-IoA)1, including AgentHub for agent registration and management as well as AgentChat for interactive communication. Furthermore, we prototype an agent discovery mechanism based on Distributed Hash Tables (DHT) and a Gossip protocol, which serves as the search engine for the internet of agentsites, enabling scalable and resilient agent registry and lookup on this new internet.
Our evaluation demonstrates that AIOS Server achieves low-latency communication, efficient task delegation, and robust coordination in agent networks. This work provides a practical foundation for building the Internet of Agentsites — a new paradigm where autonomous agents become first-class citizens of the web. The implementation is available at GitHub2.
= 1Website: https://planet.aios.foundation
[Uncaptioned image] 2Code Repository: https://github.com/agiresearch/AIOS, https://github.com/agiresearch/AIOS.Server
@@ 1 Introduction
The rapid progress of large language models (LLMs) has led to the emergence of autonomous agents capable of planning, reasoning, and interacting with humans and other agents in structured environments (Wang et al., 2024a, Yao et al., 2022, Xu et al., 2025, Wang et al., 2024b). These agents increasingly act beyond static response generation, demonstrating memory, tool usage, and long-term task execution abilities. However, most existing agent-based systems remain confined within centralized platforms, limiting openness, interoperability, and scalability in multi-agent ecosystems.
@@
This work builds on recent advances in agent system infrastructures (Mei et al., 2024, Hong et al., 2023, Deng et al., 2024), memory-enhanced LLM agents (Xu et al., 2025, Wang et al., 2024b), and agent communication protocols (Chan et al., 2023, Du et al., 2023, Liang et al., 2023). AIOS Server emphasizes modularity, interoperability, and robustness as foundational elements for open and scalable agent ecosystems. This work contributes a practical foundation for building the next-generation Internet of Agentsites, where autonomous agents operate as first-class citizens of the web, capable of decentralized collaboration, search, and interaction.
Recent works such as Chen et al. (2024) explored the idea of the internet of agents as a multi-agent framework in terms of agent team up and collaboration. However, our work is different in that we do not focus on agent team up and collaboration, but on the physical, decentralized and real-world deployment of agents across the globe (as shown in Figure 1), just like the physical Internet of Websites, where each website runs on a server at a certain location on the planet, and servers are connected through the internet to constitute the world wide web (WWW). As a result, our infrastructure is denoted as the Internet of Agentsites for differentiation and for highlighting the analogy with the Internet of Websites.
@@ 2.1 LLM-Powered Agent Systems
Recent advances in large language models (LLMs) have enabled the creation of autonomous agents that can interpret tasks, use tools, and maintain state over long interactions (Wang et al., 2024a, Yao et al., 2022, Zhou et al., 2023, Yu et al., 2023). These agents are no longer limited to single-turn responses—they can reason, remember, and act in structured workflows (Xu et al., 2025, Wang et al., 2024b). Frameworks such as Camel Li et al. (2023), OpenAGI Ge et al. (2023), AIOS Mei et al. (2024), AutoGen (Wu et al., 2023) and MetaGPT (Hong et al., 2023) define agent execution workflows and pipelines, where agents assume roles to work on task decomposition and decision-making either individually or collaboratively. Memory-augmented systems, including A-Mem and Workflow Memory (Xu et al., 2025, Wang et al., 2024b), provide persistent state tracking, improving contextual understanding and coherence. LLM-based agents now actively support applications in software development (Qian et al., 2023), web interaction (Iong et al., 2024, Deng et al., 2024), and interactive simulations (Wang et al., 2023), extending their practical utility across domains.
@@ 2.2 Agent Communication and Protocol Design
Effective coordination in agent systems requires clear, interpretable communication. Structured interaction protocols—such as ReAct (Yao et al., 2022) and intent-based dialogue models—enable agents to plan, reason, and act based on conversational or environmental context. Recent works explore debate and discussion as communication primitives to improve factuality and reasoning depth (Du et al., 2023, Liang et al., 2023, Chen et al., 2024, Pang et al., 2024, Wu et al., 2023). Message-oriented protocols like JSON-RPC are widely adopted to support structured, machine-readable interaction between agent components or subsystems. Emerging platforms such as AgentVerse (Chen et al., 2023), OS-Copilot (Wu et al., 2024), and Formal-LLM (Li et al., 2024) emphasize modular communication layers, enabling agents to interact with tools, APIs, and other services in well-defined formats.
@@ 2.3 Infrastructure and Decentralized Execution
Traditional AI agents often depend on centralized backends for registration, execution, and orchestration. This architecture limits scalability, introduces single points of failure, and reduces system adaptability. Recent efforts propose decentralized execution environments that distribute agent hosting and discovery across independent nodes. Platforms like AIOS (AI Agent Operating System) (Mei et al., 2024) highlight the need for developer-friendly, general-purpose agent development, hosting and execution frameworks. Our system builds on this foundation by introducing a runtime that integrates agent communication, registration, and discovery into a fully distributed infrastructure. Security and evaluation remain key concerns. Benchmarking frameworks such as ASB (Zhang et al., 2024) and GAIA (Mialon et al., 2023) provide standardized settings for assessing agent robustness, safety, and coordination capabilities at scale.
@@ 6 Experiments
We evaluate the AIOS server communication framework under local and cloud deployments. Experiments assess three core metrics: latency, throughput, and agent registration efficiency in a decentralized setting.
@@ 6.4.1 Global Node Distribution
Refer to caption (b) Eurasian view showing active nodes in London, Singapore, Tokyo.
@@ 7 Conclusions and Future Work
Acknowledgment
The  template used in this paper is adapted from the open-source template developed by Rongsheng Wang. We thank him for his contributions and efforts.
APPENDIX